OpenStudy (anonymous):

PLEASE HELP!! Explain, in complete sentences, the process of using Cramer’s Rule to solve the following system of equations. Provide the solution to the system. 2x + 7y = 2 3x + 8y = 2

OpenStudy (anonymous):

oh lol solve for x in one and plug it into the x in the other

OpenStudy (anonymous):

x = -2/5, y = 2/5